• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name mercado

    Log entry time 16:02:10 on September 5,2009

    Entry number 288153

    keyword=Day Shift Summary


    Day Shift Summary


    10:15 - Changed half-wave plate to IN position. Stopped Compton and HAPPEX runs for this change. Pockel Cell voltages changed to 8.532/6.167 for Gun2 and 5.025/5.000 for Gun3.

    10:22 - Resumed taking data.

    11:46 - MCC calls. Rebooting IOC. Down for 5 minutes.

    12:25 - CODA crashed. Problem with ROC25. Rebooted that crate and restarted CODA. Now complains about g0 crate, ROC31. Changed configuration to HapSimple to avoid problems.

    12:37 - Called for Controlled Access. Kent et al connect DAC signal
    from Counting House crate to Voltage channel adcx11_3

    14:15 - Kent et al out of Hall. Waiting for beam.

    14:18 - Beam is back to 100uA.

    15:29 - Beam off for like 10 minutes. Called MCC. RF cavity problem.

    15:45 - Beam is back.



    HAPPEX DAQ runs - Production at 100uA with 5.0x4.5 raster

    12019 - 12030..